A fair spinner has 9 equal sections: 4 red, 3 blue and 2 green.

It is spun twice.

What is the probability of not getting two consecutive blues?


1
Expert's answer
2021-11-08T16:13:43-0500

The probability of hitting red once is 4/9

The probability of hitting blue once is 3/9

The probability of hitting green once is 2/9

We are interested in finding the probability of not getting two consecutive blues when a spinner is spun twice.

First let us determine the probability of hitting a blue twice.

Probability of hitting a blue twice is given by 3/9*3/9=9/81

The probability of not getting two consecutive blues is 1-9/81=72/81=8/9

Therefore,  the probability of not getting two consecutive blues is 8/9.